# Exploring Multilingual Probing in Large Language Models: A Cross-Language Analysis

### v1

@@ Abstract
Probing techniques for large language models (LLMs) have primarily focused on English, overlooking the vast majority of the world’s languages. In this paper, we extend these probing methods to a multilingual context, investigating the behaviors of LLMs across diverse languages. We conduct experiments on several open-source LLM models, analyzing probing accuracy, trends across layers, and similarities between probing vectors for multiple languages. Our key findings reveal: (1) a consistent performance gap between high-resource and low-resource languages, with high-resource languages achieving significantly higher probing accuracy; (2) divergent layer-wise accuracy trends, where high-resource languages show substantial improvement in deeper layers similar to English; and (3) higher representational similarities among high-resource languages, with low-resource languages demonstrating lower similarities both among themselves and with high-resource languages. These results highlight significant disparities in LLMs’ multilingual capabilities and emphasize the need for improved modeling of low-resource languages.
@@ A Cross-Language Analysis
Daoyang Li1, Mingyu Jin2, Qingcheng Zeng3, Haiyan Zhao4, Mengnan Du4 1University of Southern California   2Rutgers University 3Northwestern University  4New Jersey Institute of Technology daoyangl@usc.edu, mengnan.du@njit.edu
@@ 2.2 Linear Classifier Probing
In our analysis, we employed linear classifier probing Ju et al. (2024a); Jin et al. (2024) to explore internal representations across various layers of LLMs. We extracted hidden states from each layer using two types of inputs (i.e., positive and negative) and utilized these features to train a logistic regression model. By evaluating the performance of this classifier, we were able to assess how well the hidden states at different layers encoded information relevant to answering factual questions or handling sentiment classification tasks. This approach provides valuable insights into the nature of the representations learned within the model.
To quantify probing differences, we employed a linear classifier approach. We define $h\in\mathbb{R}^{n\times d_{\text{model}}}$ as the hidden feature set extracted from the LLM, where $n$ is the number of samples and $d_{\text{model}}$ represents the dimensionality of the hidden layer. Each sample’s internal representation at a specific layer is denoted by $h^{(i)}\in\mathbb{R}^{1\times d_{\text{model}}}$. We utilize binary classification, assigning labels $y^{(i)}\in\{0,1\}$. The objective function for our logistic regression classifier, incorporating L2 regularization, is formulated as:
@@
$L(h^{(i)},y^{(i)};\theta)=y^{(i)}\log(\sigma(\theta^{T}h^{(i)}))+(1-y^{(i)})\log(1-\sigma(\theta^{T}h^{(i)}))$ (3)
where $\theta$ denotes the model parameters, $\lambda$ is the regularization coefficient, and $\sigma(\cdot)$ represents the sigmoid activation function. By applying this classifier to the test set, we can evaluate the LLM’s performance and gain insights into its internal representations across different languages and layers.
@@ 3 Experiment
In this section, we conduct probing experiments to compare the layer-wise accuracy and perform a correlation study to test the similarity between the probing vectors of different languages, to answer the following three research questions (RQs):
  • RQ1 - Do other languages besides English have the same probing accuracy as English?
  • RQ2 - Do other languages follow the same trends as English in different layers?
@@
  • RQ3 - Are there similarities between the probing vectors of different languages?
3.1 Experiment Settings
Models:  We utilized two open-source large language model families: Qwen Bai et al. (2023) and Gemma Team et al. (2024), to evaluate their performance and internal representations across various languages. The Qwen-1.5 models consist of 24 (0.5B & 1.8B) and 32 (7B) layers, the Gemma model has 18 (2B) and 28 (7B) layers. Different models also have different representation vector dimensions. Qwen-0.5B’s representation vector dimension is 1024, Qwen-1.8B, Gemma-2B’s representation vector dimension is 2048, Gemma-7B’s representation vector dimension is 3072, and Qwen-7B’s representation vector dimension is 4096.
Our dataset encompasses 16 languages from the whole world: English, German, French, Chinese, Spanish, Russian, Indonesian, Oriya, Hindi, Burmese, Hawaiian, Kannada, Tamil, Telugu, Kazakh, Turkmen. We categorized English, German, French, Chinese, Spanish, Russian, and Indonesian as high-resource languages, and rest of them as low-resource languages. The original language of our two datasets are English, and we used Google Translate within deep-translator python library Azam (2024) to translate them into other 15 languages, as Google Translate supports translation between over 100 languages, and achieves high accuracy compared to other translation tools.
Implementation Details:  To evaluate the performance of LLMs on each language, we use the template for the Cities dataset in English as "Judge the statement is Positive or Negative. <Statement>". The prompts of other languages utilize the same template translated by Google Translate. We applied probing techniques to assess the information encoded within each layer of these models. For our probing analysis, we selected linear classifier probing for three different experiments. The datasets are divided into a training and test set with an 8:2 ratio, and we adhered to the standard procedure for probing classifiers in LLMs, extracting feature representations from the final hidden states in the transformer at each layer of the LLMs to serve as input to the probing classifier. The linear weight parameter $\theta$ of the logistic regression classifier is regarded as the probing vector for each language and layer.
3.2 Multilingual Accuracy
In this part of our experiment, we explored (1) whether other languages besides English have the same probing accuracy as English and (2) whether they follow the same trend as English in different layers. The analysis led to two general observations. We show the results in Figure 1 and Table 1. The key findings are given as follows:
Resource Language Cities Dataset
Gemma Qwen
2B 7B 0.5B 1.8B 7B
High English 0.98 0.99 0.90 0.96 0.99
German 0.95 0.99 0.77 0.92 0.98
French 0.97 0.99 0.76 0.92 0.98
Chinese 0.69 0.76 0.70 0.75 0.88
Spanish 0.98 0.99 0.84 0.93 0.98
Russian 0.87 0.93 0.52 0.67 0.88
Indonesian 0.95 0.99 0.69 0.87 0.97
Low Oriya 0.44 0.54 0.47 0.47 0.45
Hindi 0.53 0.76 0.41 0.41 0.50
Burmese 0.60 0.81 0.33 0.37 0.44
Hawaiian 0.60 0.74 0.48 0.60 0.65
Kannada 0.60 0.72 0.43 0.42 0.40
Tamil 0.56 0.70 0.45 0.40 0.39
Telugu 0.56 0.72 0.42 0.43 0.46
Kazakh 0.66 0.75 0.43 0.44 0.67
Turkmen 0.62 0.76 0.41 0.56 0.67
3.3 Similarity Correlation of Probing Vectors
In this part of our experiment, we explored the similarities and trends between the probing vectors $\theta$ across different languages. Figure 2 and Figure 3 give an example of our results which is based on Qwen-1.8B model and Opinion dataset. We have the following observations.
  • In high-resource languages (i.e., English, German, French, Chinese, Spanish, Russian), the similarities between the probing vectors are high, but in low-resource languages, the similarities are low, even with those high-resource languages. Figure 2 is a demonstration of our observation. The similarity between German and French probing vectors is very high, but the similarity with Tamil is very low. Moreover, the probing vector similarities between Tamil and other languages are also low.
    Figure 2: Heatmap of the similarities of probing vectors correlation across languages (Model: Qwen-1.8B, Dataset: Opinion).
  • The similarity trends between high-resource languages and low-resource languages with English are completely different. Figure 3 shows that the probing vector trends of other languages exhibit similarities to English across different layers. For high-resource languages, the similarity fluctuates, whereas for low-resource languages, it remains relatively stable.
Figure 3: Cosine similarity of probing vectors with English (Model: Qwen-1.8B, Dataset: Opinion)
4 Related Work
Multilingual Abilities of LLMs. More and more researchers begin to focus on multilingual abilities of LLMs Ali and Pyysalo (2024); Jayakody and Dias (2024). Some research examines the consistency of factual knowledge across languages in multilingual pretrained language models (PLMs) Fierro and Søgaard (2022); Qi et al. (2023). Some researchers put efforts into promoting the representation of low-resource languages Abadji et al. (2022); Imani et al. (2023); Li et al. (2024). Those studies shows that LLMs still have great potential on multilingual capabilities.
5 Conclusions and Future Work
In this work, our multilingual probing experiments on LLMs reveal significant disparities in performance and representational qualities across languages. Specifically, high-resource languages consistently achieve higher probing accuracy and exhibit similar trends to English, with accuracy improving significantly in deeper layers. We also observe high similarities between probing vectors of high-resource languages, while low-resource languages demonstrate lower similarities both among themselves and with high-resource languages. These findings indicate the current limitations of LLMs in handling low-resource languages. In future, we plan to conduct research to address these gaps by developing more equitable and effective multilingual language models. Besides, we plan to extend this research to multimodal models that incorporate visual and textual information, analyzing how this affects multilingual performance.
Limitations
In this work, we use machine translation to generate the prompt templates and question sentences from English to other languages, which may introduce noise. We only experiment with 5 open-source LLMs and two datasets. In the future, we would like to expand these findings with other datasets and models to confirm how well the LLMs’ performance and representations generalize in these settings. Additionally, we just utilized linear classifier probing to do the experiments. We plan to explore more sophisticated probing methods beyond linear classifiers, which could offer deeper insights into the nature of linguistic representations within LLMs.
Acknowledgment
Appendix A Model and corresponding layers
We show the model and corresponding layers in Table 2
@@ Appendix A Model and corresponding layers
Appendix B Prompt Templates
We manually design the original natural language templates in English for both two datasets, and translate them into other languages using google translate, with a focus on expressing the exact same meaning and intent. This allows us to prevent any information differences regarding the prompt design. We present the full set of prompt templates for all languages in Figure 4.
Figure 4: Prompt templates of all languages used in experiment.
Appendix C Multilingual Accuracy
@@
In this section, we present more results on multilingual accuracy across our five evaluated models (Qwen-0.5B, Qwen-1.8B, Qwen-7B, Gemma-2B, Gemma-7B) on the Opinion dataset. We provide layer-wise accuracy plots in Figure 5 and Figure 6. These visualize how probing accuracy changes across model layers for all 16 languages. As we described in the previous section, for the Opinion dataset, the layer-wise multilingual accuracy of the five models is similar to that of the Cities dataset, which also reflects the extensiveness of our results and improves the credibility of the conclusions.
Figure 5: Additional results for multilingual accuracy of Qwen Series Model on the Opinion Dataset
Figure 6: Additional results for multilingual accuracy of Qwen Series Model on the Opinion Dataset
Appendix D Similarity Correlation of Probing Vectors
In this section, we introduce the similarities between probing vectors across different languages and show results experimental results on two datasets (Opinion, Cities) and five evaluated models (Qwen-0.5B, Qwen-1.8B, Qwen-7B, Gemma-2B, Gemma-7B) not shown in the previous section. We present:
  • Correlation heatmaps (Figure 7 - Figure 15): These visualize the pairwise similarities between probing vectors of all 16 languages. The heatmaps reveal clusters of similar languages and highlight differences between high and low-resource groups. There are slight differences in the results for different models and datasets, but overall they are still consistent with our analysis.
  • Layer-wise similarity plots (Figure 16 - Figure 24): These show how the similarity of each language’s probing vector to English changes across model layers. This reveals interesting dynamics in how different languages’ representations evolve through the model.
Figure 7: Heatmap of Similarities in Probing Vectors Correlation Across Languages Compared to English (Model: Gemma-2B, Dataset: Opinion).
Figure 8: Heatmap of Similarities in Probing Vectors Correlation Across Languages Compared to English (Model: Gemma-7B, Dataset: Opinion).
@@ Appendix D Similarity Correlation of Probing Vectors
Figure 9: Heatmap of Similarities in Probing Vectors Correlation Across Languages Compared to English (Model: Qwen-0.5B, Dataset: Opinion).
Figure 10: Heatmap of Similarities in Probing Vectors Correlation Across Languages Compared to English (Model: Qwen-7B, Dataset: Opinion).
Figure 11: Heatmap of Similarities in Probing Vectors Correlation Across Languages Compared to English (Model: Gemma-2B, Dataset: Cities).
Figure 12: Heatmap of the similarities of probing vectors correlation across languages (Model: Gemma-7B, Dataset: Cities).
Figure 13: Heatmap of the similarities of probing vectors correlation across languages (Model: Qwen-0.5B, Dataset: Cities).
Figure 14: Heatmap of the similarities of probing vectors correlation across languages (Model: Qwen-1.8B, Dataset: Cities).
Figure 15: Heatmap of the similarities of probing vectors correlation across languages (Model: Qwen-7B, Dataset: Cities).
Figure 16: Cosine Similarity of Probing Vectors with English (Model: Qwen-0.5B, Dataset: Opinion)
Figure 17: Cosine Similarity of Probing Vectors with English (Model: Qwen-7B, Dataset: Opinion)
Figure 18: Cosine Similarity of Probing Vectors with English (Model: Gemma-2B, Dataset: Opinion)
Figure 19: Cosine Similarity of Probing Vectors with English (Model: Gemma-7B, Dataset: Opinion)
Figure 20: Cosine Similarity of Probing Vectors with English (Model: Qwen-0.5B, Dataset: Cities)
Figure 21: Cosine Similarity of Probing Vectors with English (Model: Qwen-1.8B, Dataset: Cities)
Figure 22: Cosine Similarity of Probing Vectors with English (Model: Qwen-7B, Dataset: Cities)
Figure 23: Cosine Similarity of Probing Vectors with English (Model: Gemma-2B, Dataset: Cities)
Figure 24: Cosine Similarity of Probing Vectors with English (Model: Gemma-7B, Dataset: Cities)

### v2

@@ Abstract
Probing techniques for large language models (LLMs) have primarily focused on English, overlooking the vast majority of other world’s languages. In this paper, we extend these probing methods to a multilingual context, investigating the behaviors of LLMs across diverse languages. We conduct experiments on several open-source LLM models, analyzing probing accuracy, trends across layers, and similarities between probing vectors for multiple languages. Our key findings reveal: (1) a consistent performance gap between high-resource and low-resource languages, with high-resource languages achieving significantly higher probing accuracy; (2) divergent layer-wise accuracy trends, where high-resource languages show substantial improvement in deeper layers similar to English; and (3) higher representational similarities among high-resource languages, with low-resource languages demonstrating lower similarities both among themselves and with high-resource languages. These results highlight significant disparities in LLMs’ multilingual capabilities and emphasize the need for improved modeling of low-resource languages.
@@ A Cross-Language Analysis
Daoyang Li1,2, Haiyan Zhao2, Qingcheng Zeng3, Mengnan Du2 1University of Southern California, 2New Jersey Institute of Technology, 3Northwestern University daoyangl@usc.edu, mengnan.du@njit.edu
@@ 2.2 Linear Classifier Probing
In our analysis, we employed linear classifier probing Ju et al. (2024a); Jin et al. (2024) to explore internal representations across various layers of LLMs. We extracted hidden states from the residual stream of each layer using two types of inputs (i.e., positive and negative) and utilized these representations to train a logistic regression model. By evaluating the performance of trained classifier, we are able to assess how well the hidden states at different layers encoded information relevant to answering factual questions or handling sentiment classification tasks. This approach provides valuable insights into the nature of the representations learned within the model.
To perform the probing, we employed a linear classifier approach. We define $h\in{R}^{n\times d_{\text{model}}}$ as the set of hidden features extracted from the LLM, where $n$ is the number of samples and $d_{\text{model}}$ represents the dimensionality of the hidden layer. The internal representation of each sample in a specific layer is denoted by $h^{(i)}\in{R}^{1\times d_{\text{model}}}$. We utilize binary classification, assigning labels $y^{(i)}\in\{0,1\}$. The objective function for our logistic regression classifier, incorporating L2 regularization, is formulated as:
@@
$L=y^{(i)}\log(\sigma(\theta^{T}h^{(i)}))+(1-y^{(i)})\log(1-\sigma(\theta^{T}h^{(i)})),$ (3)
where $\theta$ denotes the model parameters, $\lambda$ is the regularization coefficient, and $\sigma(\cdot)$ represents the sigmoid activation function. By evaluating the accuracy of this classifier on the test set, we can evaluate the LLM’s performance and gain insights into its internal representations across different languages and layers.
@@ 3 Experiment
In this section, we conduct comprehensive probing experiments to investigate how language models process different languages. Our analysis focuses on two key aspects: comparing accuracy across different model layers and examining correlations between probing vectors of various languages. Through these experiments, we seek to answer three fundamental research questions (RQs):
  • RQ1 - Do other languages achieve comparable probing accuracy to English?
  • RQ2 - Do other languages exhibit similar layer-wise behavioral patterns to English?
@@
  • RQ3 - What are the similarities between probing vectors across different languages?
@@
3.1 Experiment Settings
In this section, we introduce the overall experimental settings of this papers.
Models:  We evaluated the performance and internal representations across various languages using two open-source LLM families: Qwen Bai et al. (2023) and Gemma Team et al. (2024). The Qwen-1.5 architecture comprises 24 layers for smaller variants (0.5B & 1.8B) and 32 layers for the larger variant (7B), while the Gemma architecture features 18 layers for the 2B model and 28 layers for the 7B model. The representation vector dimensions vary across models: 1024 for Qwen-0.5B, 2048 for both Qwen-1.8B and Gemma-2B, 3072 for Gemma-7B, and 4096 for Qwen-7B. Table 1 provides a comprehensive overview of the models and their corresponding layer configurations.
Our dataset encompasses 16 languages: English, German, French, Chinese, Spanish, Russian, Indonesian, Oriya, Hindi, Burmese, Hawaiian, Kannada, Tamil, Telugu, Kazakh, Turkmen. We categorized English, German, French, Chinese, Spanish, Russian, and Indonesian as high-resource languages, and rest of them as low-resource languages based on the volume of available digital content and linguistic resources. The original language of our two datasets are English, and we used Google Translate within deep-translator python library Azam (2024) to translate them into other 15 languages, as Google Translate supports translation between over 100 languages, and achieves high accuracy compared to other translation tools.
Model High-Resource Languages Low-Resource Languages
English German French Chinese Spanish Russian Indonesian Oriya Hindi Burmese Hawaiian Kannada Tamil Telugu Kazakh Turkmen
Gemma-2B 0.98 0.95 0.97 0.69 0.98 0.87 0.95 0.44 0.53 0.60 0.60 0.60 0.56 0.56 0.66 0.62
Gemma-7B 0.99 0.99 0.99 0.76 0.99 0.93 0.99 0.54 0.76 0.81 0.74 0.72 0.70 0.72 0.75 0.76
Qwen-0.5B 0.90 0.77 0.76 0.70 0.84 0.52 0.69 0.47 0.41 0.33 0.48 0.43 0.45 0.42 0.43 0.41
Qwen-1.8B 0.96 0.92 0.92 0.75 0.93 0.67 0.87 0.47 0.41 0.37 0.60 0.42 0.40 0.43 0.44 0.56
Qwen-7B 0.99 0.98 0.98 0.88 0.98 0.88 0.97 0.45 0.50 0.44 0.65 0.40 0.39 0.46 0.67 0.67
Implementation Details:  To evaluate the performance of LLMs on each language, we use the template for the Cities dataset in English as "Judge the statement is Positive or Negative. <Statement>". The prompts of other languages utilize the same template translated by Google Translate. This allows us to prevent any context differences regarding the prompt design. We present the full set of prompt templates for all 16 languages in Figure 3 at the Appendix. We applied probing techniques to assess the information encoded within each layer of these models. For our probing analysis, we selected linear classifier probing for our experiments. Each dataset is divided into a training and a test set with an 8:2 ratio, and we adhered to the standard procedure for probing classifiers in LLMs, extracting feature representations from the final hidden states at each layer of the LLMs to serve as input to the probing classifier. The linear weight parameter $\theta$ of the logistic regression classifier is regarded as the probing vector for each language and layer.
3.2 Multilingual Accuracy
In this section, we explored (1) whether other languages besides English have the same probing accuracy as English and (2) whether they follow the same trend as English in different layers.
@@ 3.2 Multilingual Accuracy
We present results on multilingual accuracy across our five evaluated models (Qwen-0.5B, Qwen-1.8B, Qwen-7B, Gemma-2B, Gemma-7B) on the cities and Opinion datasets. In Figure 1 and Table 2, we show the results of layer-wise probing accuracy on the Cities dataset. The results of Opinion dataset are included in Figure 4 in the Appendix. These results visualize how probing accuracy changes across model layers for all 16 languages. Based on these results, our analysis lead to two general observations as follows:
@@
Figure 2: (a) Heatmap of the similarities of probing vectors correlation across languages; (b) Cosine similarity of probing vectors with English. (Model: Qwen-1.8B, Dataset: Opinion).
3.3 Similarity Correlation of Probing Vectors
In this section, we conducted similarity analysis on probing vectors $\theta$ across languages using two visualization approaches:
  • Correlation Heatmaps: These visualize the pairwise similarities between probing vectors of all 16 languages. These highlight clustering patterns and resource-level disparities.
  • Layer-wise Similarity Plots: They measure cosine similarity between each language’s probing vector and English’s across model layers, revealing representation dynamics.
For demonstration, Figure 2 shows results from the Qwen-1.8B model and Opinion dataset. In the Appendix, we extend this analysis to all five models (Qwen-0.5B, Qwen-1.8B, Qwen-7B, Gemma-2B, Gemma-7B) and both datasets (Opinion, Cities) through Figure 5 and Figure 6. Our analysis reveals the following three key patterns:
  • The probing vectors of high-resource languages (English, German, French, Chinese, Spanish, Russian) demonstrate strong correlations with each other, as evidenced by the darker clusters in the heatmaps and consistently higher similarity curves in the trajectory plots. For instance, in the Qwen-1.8B Opinion task, German and French probing vectors maintain correlations above 0.6 with English across most layers. In contrast, low-resource languages show notably weaker correlations, both among themselves and with high-resource languages. This pattern is visible in the bright regions of the heatmaps for languages like Tamil, Telugu, and Oriya, with similarity scores typically remaining below 0.3 across all layers.
  • The evolution of similarities across model layers reveals further insights into these representational differences. High-resource languages exhibit dynamic similarity patterns with English, often peaking in middle layers before slightly decreasing, while low-resource languages maintain relatively stable, low similarity levels throughout the model layers. These patterns persist across different model sizes and architectures in both the Qwen and Gemma families, and remain consistent across the Opinion and Cities datasets.
4 Related Work
In this section, we review two lines of research that are most relevant to ours.
Multilingual Abilities of LLMs. The multilingual capabilities of LLMs have garnered increasing attention from researchers Ali and Pyysalo (2024); Jayakody and Dias (2024). Recent studies have investigated the consistency of factual knowledge across different languages in multilingual pretrained language models (PLMs) Fierro and Søgaard (2022); Qi et al. (2023). Additionally, significant efforts have been directed towards enhancing the representation of low-resource languages Abadji et al. (2022); Imani et al. (2023); Li et al. (2024). These investigations demonstrate that LLMs still possess considerable untapped potential in multilingual capabilities.
5 Conclusions and Future Work
In this work, our multilingual probing experiments on LLMs reveal significant disparities in performance and representational qualities across languages, suggesting potential limitations in how these models learn linguistic concepts. Specifically, high-resource languages consistently achieve higher probing accuracy and exhibit similar trends to English, with accuracy improving significantly in deeper layers. We also observe high similarities between probing vectors of high-resource languages, while low-resource languages demonstrate lower similarities both among themselves and with high-resource languages. These findings not only indicate the current limitations of LLMs in handling low-resource languages, but also suggest that these models may not be learning deeper linguistic concepts effectively across all languages.
In future, we plan to conduct research to address these gaps by developing more equitable and effective multilingual language models that can better capture universal linguistic concepts. Besides, we plan to extend this research to multimodal models that incorporate visual and textual information.
Limitations
In this work, we use machine translation to generate the prompt templates and question sentences from English to other languages, which may introduce noise. We only experiment with five open-source LLMs and two datasets. In the future, we would like to expand these findings with other datasets and models to confirm how well the LLMs’ performance and representations generalize in these settings. Additionally, we just utilized linear classifier probing to do the experiments. We plan to explore more sophisticated probing methods beyond linear classifiers, which could offer deeper insights into the nature of linguistic representations within LLMs.
Acknowledgment